# Eigenvalue Gaps of the Laplacian of Random Graphs

### v1

@@ Abstract.
We show that, with very high probability, the random graph Laplacian has simple spectrum. Our method provides a quantitatively effective estimate of the spectral gaps. Along the way, we establish results on affine no-gaps delocalization, no-structure delocalization, overcrowding and small entries of the eigenvectors for the Laplacian model. These findings are of independent interest.
@@ 1. Introduction
Let $G_{n}=G(n,p)$ be a random Erdős-Rényi graph with a fixed parameter $p$, and let $A_{n}$ be its adjacency matrix. Partly motivated by a question by Babai (related to the notorious graph isomorphism problem in theoretical computer science), it has been shown in [TV17, NTV17] that with high probability the spectrum sequence $\lambda_{1}(A_{n})\leq\dots\leq\lambda_{n}(A_{n})$ of $A_{n}$ is simple. More specifically
@@
[NTV17, Theorem 2.7] Let $10<p<$ 1 be fixed and let $A$ be any constant. Then for any $\delta>n^{-A}$, we have
@@
This result was extended to sparse graphs in [LV20, LL21], and to Wigner matrices (with iid subgaussian entries of mean zero, variance one) in [CJMS24] with better bounds.
From a random matrix perspective, gap sizes were in fact the original motivation for the introduction of random matrices in physics and gap statistics are a staple of the field as can be seen in [Wig55, BS10, AGZ10, EYY12, Tao13, EKYY12] and the references therein. More specifically, the minimal gap in Wigner matrices has also received much attention [Vin01, BAB13, FTW19, Bou22], where in the latest work it was shown to be universal up to some appropriate scaling.
Although these results are quite satisfactory, there remain many interesting open questions, such as to obtain the optimal quadratic repulsion probability bound $O(\delta^{2})$ (under the assumption $\delta\geq n^{{-\omega(1)}}$) in Theorem 1.1, as well as to other models of random matrices with dependent entries. The goal of this paper is to pursue the second direction for a natural model of random matrices that existing results and techniques do not seem to cover – the Laplacian matrix.
Given a graph $G_{n}$ of $n$ vertices, the (combinatorial) Laplacian of $G_{n}$, denoted by $L_{n}$, is the matrix
@@
The graph Laplacian is a discrete analogue of the Laplacian operator on manifolds. This matrix is a key object in spectral graph theory, and often of more relevance than the adjacency matrix [Nic18]. The spectrum of this matrix can reveal expansion properties, isoperimetry bounds, random walk mixing times and even the number of spanning trees of graphs [GR01]. Spectral algorithms involving the Laplacian have been spectacularly successful in fundamental tasks such as graph partitioning [HRMP95], clustering [HK92], graph drawing [FPST94], parallel computation [Sim91], graph synchronization [Ban18] and numerous others [MP93]. Laplacians arise in diverse fields ranging from physics to topological data analysis [Fyo99, WW23, BW04].
@@
For the Laplacian operator on manifolds, the eigenvalues and properties of eigenfunctions have been heavily studied in the literature, see for instance [Uhl76, CZY16, Cha84, Zho20, Wan22] and the references therein. One of the most basic questions concerns whether the eigenvalues are separate, and whether the eigenfunctions are Morse functions. Roughly speaking, it has been shown that for a generic metric, the eigenvalues are simple, and the eigenfunctions are smooth. Along these lines, genericity of simple eigenvalues for a metric graph was shown in [Fri05]. Simplicity of eigenvalues and non-vanishing of eigenfunctions of a quantum graph is considered in [BL17].
@@
For (actual) graphs, the “spectral gap" of the Laplacian matrix is the difference between the second smallest eigenvalue and the smallest eigenvalue (which is necessarily zero). This important quantity is tied to the expansion property of the graph and is of interest in many combinatorial optimization problems [Lub12]. However, there are many settings in which the other gaps between consecutive eigenvalues play an important role. In general, the mininmum gap size is intimately tied to the numerical stability of many spectral algorithms [GVL13, Dem97]. For the Laplacian in particular, the minimum gap size or the $k$-th gap size is a crucial parameter or has a specific meaning in various applications. The gap $\lambda_{n-k}-\lambda_{n-k+1}$ indicates the stability of partitioning the graph into $k$ clusters [AEGL21]. Unlike classical random walks, the mixing time for some models of quantum random walks on graphs are governed by the minimal eigenvalue gap of the Laplacian or adjacency matrix [AAKV01, Kem03, CLR20b]. These quantum walks are the basis of efficient quantum algorithms that far outperform classical algorithms and many of these quantum algorithms implicitly assume that the spectrum of the Laplacian matrix for the graphs are simple [AAKV01, CLR20a, CCD+03, Chi04]. For many graph-based learning tasks, graph neural networks (GNNs) are the industry standard solution method [ZCH+20]. Graphs with many symmetries, however, often lead to slow convergence rates and low accuracy [LL22]. One promising solution to this is to endow the vertices with more features to break the symmetries, often by embedding the graph in Euclidean space using the Laplacian matrix [ACGL20, LWWL20]. In this algorithm, the rate of convergence is specified by the size of the smallest eigenvalue gap of the Laplacian matrix [WYZL22, ZL20, MWW23, GBR20]. Some further applications of the minimal eigenvalue gaps are found in graph synchronization [Nyb14, NGB15].
Motivated by all these directions, we show the following
Theorem 1.2 (Eigenvalue gaps of Laplacian, main result).
@@
In particular, the spectrum is simple with very high probability. This provides theoretical justification for the empirical success or implicit assumptions in the cited applications. More quantitatively, by taking the union bound over all $i$-th gaps, with probability $1-o(1)$ we see that the minimum gap in fact has order at least $n^{{-3/2}-o(1)}$.
Notations. Throughout this paper, we regard $n$ as an asymptotic parameter going to infinity (in particular, we will implicitly assume that $n$ is larger than any fixed constant, as our claims are all trivial for fixed $n$), and allow all mathematical objects in the paper to implicitly depend on $n$ unless they are explicitly declared to be “fixed” or “constant”. We write $X=O(Y),X\ll Y$, or $Y=\Omega(X),Y\gg X$ to denote the claim that $|X|\leq CY$ for some fixed constant $C$. We write $X=\Theta(Y)$ if $X\gg Y$ and $Y\gg X$. We also use $o(1)$ (and $\omega(1)$ respectively) to denote positive quantities that tend to zero (infinity) as $n\to\infty$.
For a square matrix $X$ of size $n$ and a number $\lambda$, for short we will write $X-\lambda$ instead of $X-\lambda I_{n}$. We denote ${\mathbf{r}}_{i}(X)$ and ${\mathbf{c}}_{i}(X)$ by the $i$-th row and column of $X$. For historical reason, for an $n\times k$ or $k\times n$ matrix $X$ with $k\leq n$, we let $0\leq\sigma_{k}(X)\leq\cdots\leq\sigma_{1}(X)$ be the singular values of $X$, while reversely, for an $n\times n$ symmetric matrix $X$, we let $\lambda_{1}(X)\leq\cdots\leq\lambda_{n}(X)$ denote its eigenvalues and $\delta_{i}(X)=\lambda_{i+1}(X)-\lambda_{i}(X)$ be its $i^{\text{th}}$ eigenvalue gap.
We denote the discrete interval $[m]=\left\{1,2,\cdots,m\right\}$. For ${\mathbf{v}}\in{\mathbb{R}}^{n}$ and $I\subset[n]$, we denote ${\mathbf{v}}_{I}\in{\mathbb{R}}^{I}$ the vector with entries $v_{i},\text{ for }i\in I$. Finally, all the norms in this note, if not specified, will be the usual $l_{2}$-norm.
2. Other main results and our approach
@@ 2.1. Eigenvalue gaps for the centered model
It is also natural to study the centered Laplacian (which may go by other names such as Laplacian matrix or Laplacian-type matrix)
@@
In a more general setting, we can start from a Wigner matrix $X_{n}=(x_{ij})_{1\leq i,j\leq n}$ (with iid entries $x_{ij},1\leq i<j\leq n$ of mean zero and variance one) and form a centered Laplacian $\overline{X}_{n}$ from it by letting $\overline{x}_{ii}=\sum_{j,j\neq i}x_{ij}$ and $\overline{x}_{ij}=-x_{ij}$. Hence our matrix $\overline{L}_{n}/\sqrt{p(1-p)}$ belongs to this family of random matrices.
It is known (see [DJ10]) that the empirical spectral distribution of a random centered Laplacian converges weakly to the free convolution of the semicircular law and the standard gaussian, $\rho_{sc}\boxplus{\mathbf{N}}(0,1)$. The gaps between extreme eigenvalues of the Laplacian have been studied [Jia12, CO07, CLO+22]. Furthermore, among other things, it was shown in [HL20] that individual eigenvalue gaps in the bulk of the spectrum converge to the gap behavior of the GOE. However, these asymptotic results do not cover the basic question of whether $\overline{L}_{n}$ has simple spectrum with high probability. Here we show
@@
We next outline an approach  [^1] There are more advanced approaches to this problem, see for instance [TV11, Eqn. 36]. from [NTV17] to prove Theorem 1.1, and single out the difficulty of the current model. Assume that $X_{n}$ is a symmetric/Hermitian matrix where $\lambda_{i+1}-\lambda_{i}\leq\delta/n^{1/2}$.
@@
$X_{n}=\begin{pmatrix}X_{n-1}&{\mathbf{c}}_{n}\\ {\mathbf{c}}_{n}^{*}&x_{nn}\end{pmatrix}$ (1)
where ${\mathbf{c}}_{n}$ is the last column vector. From the Cauchy interlacing law, we observe that $\lambda_{i}(X_{n})\leq\lambda_{i}(X_{n-1})\leq\lambda_{i+1}(X_{n})$. Let ${\mathbf{u}}$ be the (unit) eigenvector of $\lambda_{i}(X_{n})$; we write ${\mathbf{u}}=({\mathbf{w}},b)$, where ${\mathbf{w}}$ is a vector of length $n-1$ and $b$ is a scalar. We have
@@
Extracting the top $n-1$ components of this equation we obtain
@@
Let ${\mathbf{v}}$ be the unit eigenvector of $X_{n-1}$ corresponding to $\lambda_{i}(X_{n-1})$. By multiplying with ${\mathbf{v}}^{T}$, we obtain
@@
We conclude that, if $\lambda_{i+1}-\lambda_{i}\leq\delta n^{-1/2}$ holds, then $|b{\mathbf{v}}^{T}{\mathbf{c}}_{n}|\leq\delta n^{-1/2}$. If we assume that $|b|\gg n^{-1/2+o(1)}$, then
@@
Thus we have reduced the problem to bounding the probability with respect to $X_{n-1}$ and ${\mathbf{c}}_{n}$ that $X_{n-1}$ has an eigenvector ${\mathbf{v}}$ and such that $|{\mathbf{v}}^{T}{\mathbf{c}}_{n}|=O(n^{o(1)}\delta)$. The plan is then divided into two separate steps:
@@
The structure of the Laplacian matrix $L_{n}$ (and $\overline{L}_{n}$) pose several challenges when compared with that of the symmetric/Hermitian matrix $A_{n}$ (which plays the role of $X_{n}$ above). The most apparent difficulty is the dependence among the entries above the diagonal. For instance if we fix all the entries of $L_{n-1}$, the upper left principle minor of size $n-1$ of $L_{n}$, then $L_{n}$ is determined. In other words, even if one can execute Step 1 for $L_{n-1}$, we cannot rely on Step 2 because ${\mathbf{c}}_{n}(L_{n})$ is already determined after fixing $L_{n-1}$. To avoid this deadlock, we will proceed as follows (using the model $G(n,1/2)$ as an example – the approach naturally extends to $G(n,p)$ and beyond). First, assume that the vertices of $G=G_{n}$ are ordered as $(v_{1},\dots,v_{n})$. We sample the entries $a_{ij}$ of the adjacency matrix (via $G(n,1/2)$) and compute the degrees $d_{1},\dots,d_{n}$ of the vertices, and then form $\overline{L}_{n}$. Let $\overline{L}_{n}^{\prime}$ be the matrix obtained from $L_{n}$ by fixing all rows and columns, except that the last two columns ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ (as well as the last two rows ${\mathbf{r}}_{n-1},{\mathbf{r}}_{n}$) are replaced by ${\mathbf{c}}_{n-1}+{\mathbf{c}}_{n}$ and ${\mathbf{c}}_{n}-{\mathbf{c}}_{n-1}$  [^2] In fact, to preserve the spectrum, we will replace ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ by $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $({\mathbf{c}}_{n}-{\mathbf{c}}_{n-1})/\sqrt{2}$ (see Section 7), but let’s ignore this minor problem for now as the proofs remain the same..
@@
Next, we reshuffle the neighbors of $v_{n-1}$ and $v_{n}$ as follows: consider the set $I_{n-2}$ of indices $21\leq j\leq n-$ 2 where $v_{j}$ is connected to exactly one of $v_{n-1}$ or $v_{n}$. Then for each $j\in I_{n-2}$, we flip a fair coin to either keep or swap whether each of $(v_{j},v_{n-1})$ and $(v_{j},v_{n})$ is an edge (see Figure 1). In other words, if $X=(x_{1},\dots,x_{n-2})$ and $Y=(y_{1},\dots,y_{n-2})$ are the (restricted to the first $n-2$ coordinates) column vectors of $A_{n}$ (or $L_{n},\overline{L}_{n}$) associated to $v_{n-1}$ and $v_{n}$, then $I_{n-2}$ is the collection of indices $j$ where $(x_{j},y_{j})=(0,1)$ or $(1,0)$ (respectively $(0,-1)$ or $(-1,0)$ in the $L_{n}$ case, and $(1/2,-1/2)$ or $(-1/2,1/2)$ in the $\overline{L}_{n}$ case). We then flip a fair coin to reassign $(x_{j},y_{j})$ to $(0,1)$ or $(1,0)$. To some extent, this reshuffling is similar to [McK81] and [Coo17, LLT+21, NW22, NW24] where shufflings/switchings were used within random graphs and random matrices. However our implementation here is rather straightforward. We can easily see that the above process does not change the law of $\overline{L}_{n}$. In terms of $\overline{L}_{n}^{\prime}$, this process searches for the index set $I_{n-2}\subset[n-2]$ where the entries of ${\mathbf{c}}_{n-1}(\overline{L}^{\prime}_{n})$ (or the entries of ${\mathbf{c}}_{n-1}(\overline{L}_{n})+{\mathbf{c}}_{n}(\overline{L}_{n})$) is zero, and then for each $j\in I_{n-2}$, the entries of ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ take values ±$\pm 1$ (independently from each other, and from $\overline{L}^{\prime}_{n-1}$) with probability 1/2. By Chernoff’s bound, with high probability $I_{n-2}$ has size approximately $(n-2)/2$ (or $(n-2)p$ in general). With the new randomness of ${\mathbf{c}}_{n}(\overline{L}_{n}^{\prime})_{|_{I_{n-2}}}$, we can now move to Step 2 to study (3), which can be rewritten as
@@
While we can proceed with Step 2, a new problem arises, that in Step 1 it is not enough to know that ${\mathbf{v}}$ does not have structure, but we have to show that ${\mathbf{v}}_{I}$ does not have structure for any eigenvector of $\overline{L}_{n-1}^{\prime}$. As far as we are concerned, the property of eigenvectors not having local structure has not been studied before in the literature, except in the no-gaps delocalization aspect. (Although in the finite field setting, this local aspect has been studied in [NW22] for null vectors.)

### v2

@@ Abstract.
We show that, with very high probability, the random graph Laplacian has simple spectrum. Our method provides a quantitatively effective estimate of the spectral gaps. Along the way, we establish results on affine no-gaps delocalization, no-structure delocalization, overcrowding and the presence of small entries in the eigenvectors of the Laplacian model. These findings are of independent interest.
Key words and phrases:
Eigenvalue Gaps, Spectral Gaps, Graph Laplacian, Quantum Random Walks, Random Matrices, Delocalization of Eigenvectors
2020 Mathematics Subject Classification:
60B20, 60C05, 05C80, 15B52
K. Luh has been partially supported by Simons Grant MP-TSM-00001988.
@@ 1. Introduction
Let $G_{n}=G(n,p)$ be a random Erdős-Rényi graph with a fixed parameter $p$, and let $A_{n}$ be its adjacency matrix. Partly motivated by a question posed by Babai (related to the well-known graph isomorphism problem in theoretical computer science), it has been shown in [64, 52] that with high probability, the spectrum $\lambda_{1}(A_{n})\leq\dots\leq\lambda_{n}(A_{n})$ of $A_{n}$ is simple. More precisely,
@@
[52, Theorem 2.7] Let $10<p<$ 1 be fixed, and let $A$ be any constant. For any $\delta>n^{-A}$, we have
@@
This result was extended to sparse graphs in [47, 45], and to Wigner matrices (with i.i.d. subgaussian entries of mean zero and variance one) in [13] with better bounds.
From a random matrix perspective, gap sizes were, in fact, the original motivation for introducing random matrices in physics. Further, gap statistics are a staple of the field as can be seen in [71, 5, 3, 29, 62, 28] and the references therein. More specifically, the minimal gap in Wigner matrices has also received much attention [67, 8, 30, 11], where, in the latest work, it was shown to be universal up to some appropriate scaling.
Although these results are quite satisfactory, many interesting open questions remain. These include obtaining the optimal quadratic repulsion probability bound $O(\delta^{2})$ (under the assumption $\delta\geq n^{{-\omega(1)}}$) in Theorem 1.1, as well as extending results to other models of random matrices with dependent entries. The goal of this paper is to pursue the second direction for a natural model of random matrices that existing results and techniques do not seem to cover – the Laplacian matrix.
Given a graph $G_{n}$ with $n$ vertices, the (combinatorial) Laplacian of $G_{n}$, denoted by $L_{n}$, is the matrix
@@
The graph Laplacian is a discrete analogue of the Laplacian operator on manifolds. This matrix is a key object in spectral graph theory and is often more relevant than the adjacency matrix. The spectrum of this matrix can reveal expansion properties, isoperimetric bounds, random walk mixing times, and even the number of spanning trees of a graph [35]. Spectral algorithms involving the Laplacian have been spectacularly successful in fundamental tasks such as graph partitioning [38], clustering [37], graph drawing [31], parallel computation [61], graph synchronization [6], and numerous others [50]. Laplacians arise in diverse fields ranging from physics to topological data analysis [33, 70, 7].
@@
For the Laplacian operator on manifolds, the eigenvalues and properties of eigenfunctions have been heavily studied in the literature; see, for instance, [65, 20, 19, 73, 69] and the references therein. One of the most basic questions concerns whether the eigenvalues are distinct and whether the eigenfunctions are Morse functions. Roughly speaking, it has been shown that for a generic metric, the eigenvalues are simple and the eigenfunctions are smooth. Along these lines, genericity of simple eigenvalues for a metric graph was shown in [32]. Simplicity of eigenvalues and non-vanishing of eigenfunctions of a quantum graph are considered in [9].
@@
For (actual) graphs, the “spectral gap” of the Laplacian matrix is the difference between the second smallest eigenvalue and the smallest eigenvalue (which is necessarily zero). This important quantity is tied to the expansion property of the graph and is of interest in many combinatorial optimization problems [46]. However, there are many settings in which the gaps between other consecutive eigenvalues play an important role. In general, the minimum gap size is closely tied to the numerical stability of many spectral algorithms [36, 26]. For the Laplacian in particular, the minimum gap size or the $k$-th gap size is a crucial parameter and has a specific meaning in various applications. The gap $\lambda_{n-k}-\lambda_{n-k+1}$ indicates the stability of partitioning the graph into $k$ clusters [4].
Next, we highlight a key motivation for our work: quantum random walks on graphs. Unlike classical random walks, where mixing times are governed by the spectral gap of the transition matrix, the mixing time of quantum random walks is determined by the smallest eigenvalue gap of the Laplacian or the adjacency matrix [2, 41, 18]. These quantum walks serve as the foundation for efficient quantum algorithms that often significantly outperform their classical counterparts. Many of these quantum algorithms implicitly assume that the spectrum of the graph Laplacian is simple [2, 17, 21, 22]. More specifically, let $G$ be a graph with vertex set $V=\{1,\dots,n\}$. The Hamiltonian governing the quantum walk on $G$ is given by $\gamma L_{G}$, where $L_{G}$ is the graph Laplacian and $\gamma$ is a scaling factor. Given an initial state $|\psi_{0}\rangle\in\mathcal{H}$, the state of the walker at time $t$ evolves as $|\psi(t)\rangle=e^{-i\gamma L_{G}t}|\psi_{0}\rangle$. The probability that the walker is at node $|f\rangle$ after time $T$ is
[…]
The limiting distribution is given by $P_{f}(T\to\infty):=\lim_{T\to\infty}P_{f}(T)=\sum_{i=1}^{n}|\langle f|{\mathbf{v}}_{i}\rangle\langle{\mathbf{v}}_{i}|\psi_{0}\rangle|^{2}$, where ${\mathbf{v}}_{i}$ are the eigenvectors of $L_{G}$. It can be shown that the discrepancy in the $L^{1}$-norm is
$D(P_{T})=\sum_{f}\left|P_{f}(T)-P_{f}(T\to\infty)\right|\leq\sum_{i\neq l}\frac{2|\langle{\mathbf{v}}_{i}|\psi_{0}\rangle|\cdot|\langle{\mathbf{v}}_{l}|\psi_{0}\rangle|}{T|\lambda_{i}-\lambda_{l}|}.$
Thus, controlling the sum $\sum_{i\neq l}\frac{1}{|\lambda_{i}-\lambda_{l}|}$ is crucial, as it quantifies the separation of the entire spectrum.
Beyond the applications above, the minimal eigenvalue gap of the Laplacian matrix also plays a significant role in many graph-based learning tasks. In particular, graph neural networks (GNNs) have become the industry standard for graph-based machine learning tasks [72]. However, highly symmetric graphs often lead to slow convergence rates and low accuracy [42]. A promising solution is to enrich vertex features to break symmetries, often by embedding the graph into Euclidean space using the Laplacian matrix [1, 43]. In these algorithms, the rate of convergence is determined by the size of the smallest eigenvalue gap of the Laplacian matrix [68, 74, 48, 34]. Further applications of minimal eigenvalue gaps can be found in graph synchronization [56, 57].
Motivated by all these directions, we show the following main result.
Theorem 1.2 (Eigenvalue Gaps of Laplacian).
@@
In particular, the spectrum is simple with high probability. This provides a theoretical justification for the empirical success or implicit assumptions in the cited applications. More quantitatively, by taking the union bound over all $i$-th gaps, we find that, with probability $1-o(1)$, the minimum gap is at least of order $n^{{-3/2}-o(1)}$.
Notations. Throughout this paper, we regard $n$ as an asymptotic parameter tending to infinity (in particular, we implicitly assume that $n$ is larger than any fixed constant, as our claims are all trivial for fixed $n$), and allow all mathematical objects in the paper to implicitly depend on $n$ unless they are explicitly declared to be “fixed” or “constant”. We write $X=O(Y),X\ll Y$, or $Y=\Omega(X),Y\gg X$ to denote the claim that $|X|\leq CY$ for some fixed constant $C$. We write $X=\Theta(Y)$ if $X\gg Y$ and $Y\gg X$. We also use $o(1)$ (and $\omega(1)$ respectively) to denote positive quantities that tend to zero (infinity) as $n\to\infty$.
For a square matrix $X$ of size $n$ and a number $\lambda$, for brevity, we write $X-\lambda$ instead of $X-\lambda I_{n}$. We use ${\mathbf{r}}_{i}(X)$ and ${\mathbf{c}}_{i}(X)$ to denote the $i$-th row and column of $X$, respectively. For historical reasons, for an $n\times k$ or $k\times n$ matrix $X$ with $k\leq n$, we let $0\leq\sigma_{k}(X)\leq\cdots\leq\sigma_{1}(X)$ be the singular values of $X$. Conversely, for an $n\times n$ symmetric matrix $X$, we let $\lambda_{1}(X)\leq\cdots\leq\lambda_{n}(X)$ denote its eigenvalues and $\delta_{i}(X)=\lambda_{i+1}(X)-\lambda_{i}(X)$ be its $i^{\text{th}}$ eigenvalue gap.
We denote the discrete interval $[m]=\left\{1,2,\cdots,m\right\}$. For ${\mathbf{v}}\in{\mathbb{R}}^{n}$ and $I\subset[n]$, we denote ${\mathbf{v}}_{I}\in{\mathbb{R}}^{I}$ the vector with entries $v_{i},\text{ for }i\in I$. Finally, throughout this paper, if not specified, any norm that appears is the standard Euclidean norm.
2. Additional Main Results and Our Approach
@@ 2.1. Eigenvalue Gaps for the Centered Model
It is also natural to study the centered Laplacian (also sometimes referred to as the Laplacian matrix or a Laplacian-type matrix)
@@
In a more general setting, we can start from a Wigner matrix $X_{n}=(x_{ij})_{1\leq i,j\leq n}$ (with i.i.d. entries $x_{ij},1\leq i<j\leq n$ of mean zero and variance one) and form a centered Laplacian $\overline{X}_{n}$ from it by letting $\overline{x}_{ii}=\sum_{j,j\neq i}x_{ij}$ and $\overline{x}_{ij}=-x_{ij}$. Thus, the matrix $\overline{L}_{n}/\sqrt{p(1-p)}$ belongs to this family of random matrices.
It is known (see [27]) that the empirical spectral distribution of a random centered Laplacian converges weakly to the free convolution of the semicircular law and the standard Gaussian, $\rho_{sc}\boxplus{\mathbf{N}}(0,1)$. The gaps between extreme eigenvalues of the Laplacian have been studied [40, 23, 12]. Furthermore, among other things, it was shown in [39] that individual eigenvalue gaps in the bulk of the spectrum converge to the gap behavior of the GOE. However, these asymptotic results do not cover the basic question of whether $\overline{L}_{n}$ has simple spectrum with high probability. In this paper, we show the following result.
@@
Next, we outline an approach  [^1] There are more advanced approaches to this problem, see for instance [63, Eqn. 36]. from [52] to prove Theorem 1.1, and highlight the challenges in analyzing the current model. Assume that $X_{n}$ is a symmetric/Hermitian matrix where $\lambda_{i+1}-\lambda_{i}\leq\delta/n^{1/2}$.
@@
$X_{n}=\begin{pmatrix}X_{n-1}&{\mathbf{c}}_{n}\\ {\mathbf{c}}_{n}^{*}&x_{nn}\end{pmatrix},$ (1)
where ${\mathbf{c}}_{n}$ is the last column of $X_{n}$ (without the final entry). From the Cauchy interlacing law, we observe that $\lambda_{i}(X_{n})\leq\lambda_{i}(X_{n-1})\leq\lambda_{i+1}(X_{n})$. Let ${\mathbf{u}}$ be a (unit) eigenvector of $X_{n}$ with associated eigenvalue $\lambda_{i}(X_{n})$. We write ${\mathbf{u}}=({\mathbf{w}},b)$, where ${\mathbf{w}}$ is a vector of length $n-1$ and $b$ is a scalar. We have
@@
Extracting the top $n-1$ components of this equation, we obtain
@@
Let ${\mathbf{v}}$ be a unit eigenvector of $X_{n-1}$ corresponding to $\lambda_{i}(X_{n-1})$. By multiplying the above equation on the left by ${\mathbf{v}}^{T}$, we get that
@@
We conclude that if $\lambda_{i+1}-\lambda_{i}\leq\delta n^{-1/2}$ holds, then $|b{\mathbf{v}}^{T}{\mathbf{c}}_{n}|\leq\delta n^{-1/2}$. If we assume that $|b|\gg n^{-1/2+o(1)}$, then
@@
Thus, we have reduced the problem to bounding the probability with respect to $X_{n-1}$ and ${\mathbf{c}}_{n}$ that $X_{n-1}$ has an eigenvector ${\mathbf{v}}$ such that $|{\mathbf{v}}^{T}{\mathbf{c}}_{n}|=O(n^{o(1)}\delta)$. The plan is then divided into two separate steps:
@@
The structure of the Laplacian matrix $L_{n}$ (and $\overline{L}_{n}$) pose several challenges when compared with that of the symmetric/Hermitian matrix $A_{n}$ (which plays the role of $X_{n}$ above). The most apparent difficulty is the dependence among the entries above the diagonal. For instance if we fix all the entries of $L_{n-1}$, the upper left principal minor of size $n-1$ of $L_{n}$, then $L_{n}$ is determined. In other words, even if one can execute Step 1 for $L_{n-1}$, we cannot rely on Step 2 because ${\mathbf{c}}_{n}(L_{n})$ is already determined after fixing $L_{n-1}$. To avoid this deadlock, we will proceed as follows (using the model $G(n,1/2)$ as an example – the approach naturally extends to $G(n,p)$ and beyond). First, assume that the vertices of $G=G_{n}$ are ordered as $(v_{1},\dots,v_{n})$. We sample the entries $a_{ij}$ of the adjacency matrix (via $G(n,1/2)$) and compute the degrees $d_{1},\dots,d_{n}$ of the vertices and then form $\overline{L}_{n}$. Let $\overline{L}_{n}^{\prime}$ be the matrix obtained from $L_{n}$ by fixing all rows and columns, except that the last two columns ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ (as well as the last two rows ${\mathbf{r}}_{n-1},{\mathbf{r}}_{n}$) are replaced by ${\mathbf{c}}_{n-1}+{\mathbf{c}}_{n}$ and ${\mathbf{c}}_{n}-{\mathbf{c}}_{n-1}$  [^2] In fact, to preserve the spectrum, we will replace ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n}$ by $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $({\mathbf{c}}_{n}-{\mathbf{c}}_{n-1})/\sqrt{2}$ (see Section 7), but let’s ignore this minor problem for now as the proofs remain the same..
Next, we reshuffle the neighbors of $v_{n-1}$ and $v_{n}$ as follows: consider the set $I_{n-2}$ of indices $21\leq j\leq n-$ 2 where $v_{j}$ is connected to exactly one of $v_{n-1}$ or $v_{n}$. Then for each $j\in I_{n-2}$, we flip a fair coin to either keep or swap whether each of $(v_{j},v_{n-1})$ and $(v_{j},v_{n})$ is an edge (see Figure 1).
@@
In other words, if $X=(x_{1},\dots,x_{n-2})$ and $Y=(y_{1},\dots,y_{n-2})$ are the (restricted to the first $n-2$ coordinates) column vectors of $A_{n}$ (or $L_{n},\overline{L}_{n}$) associated to $v_{n-1}$ and $v_{n}$, then $I_{n-2}$ is the collection of indices $j$ where $(x_{j},y_{j})=(0,1)$ or $(1,0)$ (respectively $(0,-1)$ or $(-1,0)$ in the $L_{n}$ case, and $(1/2,-1/2)$ or $(-1/2,1/2)$ in the $\overline{L}_{n}$ case). We then flip a fair coin to reassign $(x_{j},y_{j})$ to $(0,1)$ or $(1,0)$. To some extent, this reshuffling is similar to [49] and [24, 44, 54, 55] where shufflings/switchings were used within random graphs and random matrices. However, our implementation here is rather straightforward. We can easily see that the above process does not change the law of $\overline{L}_{n}$. In terms of $\overline{L}_{n}^{\prime}$, this process searches for the index set $I_{n-2}\subset[n-2]$ where the entries of ${\mathbf{c}}_{n-1}(\overline{L}^{\prime}_{n})$ (or the entries of ${\mathbf{c}}_{n-1}(\overline{L}_{n})+{\mathbf{c}}_{n}(\overline{L}_{n})$) are zero, and then for each $j\in I_{n-2}$, the entries of ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ take values ±$\pm 1$ (independently from each other, and from $\overline{L}^{\prime}_{n-1}$) with probability 1/2. By Chernoff’s bound, with high probability, $I_{n-2}$ has size approximately $(n-2)/2$ (or $(n-2)p$ in general). With the new randomness of ${\mathbf{c}}_{n}(\overline{L}_{n}^{\prime})_{|_{I_{n-2}}}$, we can now move to Step 2 to study (3), which can be rewritten as
@@
While we can proceed with Step 2, a new problem arises: in Step 1 it is not enough to know that ${\mathbf{v}}$ does not have structure, but we have to show that ${\mathbf{v}}_{I}$ does not have structure for any eigenvector of $\overline{L}_{n-1}^{\prime}$. To the authors’ best knowledge, the property of eigenvectors not having local structure has not been studied before in the literature, except in the no-gaps delocalization aspect. (Although in the finite field setting, this local aspect has been studied in [54] for null vectors.)